(D² + 3DD '+ 2D'²) z = x+y
1
Expert's answer
2019-08-26T05:00:11-0400

SOLUTION

Let us use some known facts. Let the given differential equation be


"F(D,D')=f(x,y)"

Factorize "F(D,D')" into linear factors. Then use the following results:

 Corresponding to each non-repeated factor "\left(bD-aD'-c\right)" , the part of C.F. is taken as


"\exp\left(\frac{cx}{b}\right)\cdot\varphi(by+ax),\quad if\quad b\neq0"

In our case,


"\left(D^2+3DD'+2D'^2\right)z=x+y\longrightarrow z(x,y)=C.F.+P.I."

0 STEP: We factor the expression

"\left(D^2+3DD'+2D'^2\right)=\left(D^2+DD'\right)+\left(2DD'+2D'^2\right)="

"=D\left(D+D'\right)+2D'\left(D+D'\right)=\left(D+D'\right)\left(D+2D'\right)"

Conclusion,


"\boxed{\left(D^2+3DD'+2D'^2\right)z=\left(D+D'\right)\left(D+2D'\right)z}"

1 STEP: Let find C.F.


"\left\{\begin{array}{c}
\left(D+D'\right)z\\
\left(bD-aD'-c\right)z
\end{array}\right.\rightarrow
\left\{\begin{array}{c}
a=-1\\
b=1\\
c=0
\end{array}\right.\rightarrow"

"(C.F.)_1=\exp\left(\frac{0\cdot x}{1}\right)\cdot\varphi(1\cdot y+(-1)x)"



"\boxed{(C.F.)_1=\varphi_1(y-x),\quad where\,\,\,\varphi_1\,\,\,is \,\,\,arbitrary\,\,\,function}"

"\left\{\begin{array}{c}
\left(D+2D'\right)z\\
\left(bD-aD'-c\right)z
\end{array}\right.\rightarrow
\left\{\begin{array}{c}
a=-2\\
b=1\\
c=0
\end{array}\right.\rightarrow"


"(C.F.)_2=\exp\left(\frac{0\cdot x}{1}\right)\cdot\varphi(1\cdot y+(-2)x)"


"\boxed{(C.F.)_2=\varphi_2(y-2x),\quad where\,\,\,\varphi_2\,\,\,is \,\,\,arbitrary\,\,\,function}"

Then,


"C.F.=(C.F.)_1+(C.F.)_2=\varphi_1(y-x)+\varphi_2(y-2x)"

2 STEP: Let find P.I.


"P.I.=\frac{1}{(D+D')(D+2D')}(x+y)=\frac{1}{D\left(1+\frac{D'}{D}\right)D\left(1+\frac{2D'}{D}\right)}(x+y)="

"=\frac{1}{D^2}\left(1+\frac{D'}{D}\right)^{-1}\left(1+\frac{2D'}{D}\right)^{-1}(x+y)="

"=\frac{1}{D^2}\left[1-\frac{D'}{D}+\ldots\right]\left[1-\frac{2D'}{D}+\ldots\right](x+y)="


"=\left[\frac{1}{1+x}=1-x+x^2-x^3+\ldots,|x|<1\right]="

"=\frac{1}{D^2}\left[1-\frac{D'}{D}-\frac{2D'}{D}+\frac{2D'^2}{D^2}+\ldots\right](x+y)="


"=\frac{1}{D^2}\left[1-\frac{3D'}{D}+\frac{2D'^2}{D^2}+\ldots\right](x+y)="

"=\frac{1}{D^2}\left[(x+y)-\frac{3}{D}\left(\frac{\partial}{\partial y}(x+y)\right)+\frac{2}{D^2}\left(\frac{\partial^2}{\partial^2 y}(x+y)\right)+\ldots\right]="

"=\frac{1}{D^2}\left[(x+y)-\frac{3}{D}\left(1\right)+\frac{2}{D^2}\left(0\right)+\ldots\right]="


"=\frac{1}{D^2}\left[(x+y)-\frac{3}{D}\left(1\right)\right]=\left(\frac{1}{D}\equiv\int dx\right)="

"=\frac{1}{D^2}\left[(x+y)-3\int1dx\right]=\frac{1}{D^2}\left[(x+y)-3x\right]="


"=\frac{1}{D}\left[\int(y-2x)dx\right]=\int\left(xy-x^2\right)dx=\frac{x^2y}{2}-\frac{x^3}{3}"

"\boxed{P.I.=\frac{x^2y}{2}-\frac{x^3}{3}}"

Conclusion,


"z(x,y)=C.F.+P.I.=\varphi_1(y-x)+\varphi_2(y-2x)+\frac{x^2y}{2}-\frac{x^3}{3}"

ANSWER

"\left\{\begin{array}{c}
z(x,y)=\varphi_1(y-x)+\varphi_2(y-2x)+\displaystyle\frac{x^2y}{2}-\frac{x^3}{3}\\[0.4cm]
where\,\,\,\varphi_1\,\,\,and\,\,\,\varphi_2\,\,\,are\,\,\,arbitrary\,\,\,functions
\end{array}\right."